# Geometry of orofacial neuromuscular signals: speech articulation decoding using surface electromyography

### v1

@@ 5 Methods and data properties
With sEMG articulations represented on the Riemannian manifold, we can apply methods described in Huang and Van Gool (2017) for SPD matrix learning to decode different articulations. We describe the architecture of the neural network in figure 14 and briefly recapitulate the core formulations in Huang and Van Gool (2017) here.
@@ 6.1 Experiment 1
Word and phonetic articulations in table LABEL:tab:experimentPart1 are produced in 4 different sessions. Subjects articulated phonemes and words in an audible manner in the first 2 sessions and in a silent manner in the next 2 sessions. Within each session, subjects articulated a given word or phoneme 5 times. Therefore, we have 10 instances each of audible and silent articulations for all words and phonemes. We use first 3 instances from each session for training and last two instances for testing the models. Therefore, we have 6 instances of each word or phoneme for training and 4 instances for testing. Phonetic articulation data of subject 11 was corrupted during acquisition and is unavailable. For MDM and architecture in figure 14, edge matrices are created using the time-context window of 1.5s (the entire duration of articulation). For GRU architecture in figure 15, edge matrices are created with a sliding window of context size of 150ms and step size of 30ms. Results for the experiments in table LABEL:tab:experimentPart1 are given here. The presented accuracy values are averaged over 10 random seeds (the model in figure 14 is trained for 1000 epochs and the model in figure 15 is trained for 150 epochs. Best accuracy on the test set in reported). Model in figure 14 contains about 5000 to 11000 parameters (depending on the size of the linear classification head) and the model in figure 15 has approximately 150000 parameters.
@@ Acknowledgments
This work was supported by awards to LMM from: Accenture, through the Accenture Labs Digital Experiences group; CITRIS and the Banatao Institute at the University of California; the University of California Davis School of Medicine (Cultivating Team Science Award); the University of California Davis Academic Senate; a UC Davis Science Translation and Innovative Research (STAIR) Grant; and the Child Family Fund for the Center for Mind and Brain.
HTG is supported by Neuralstorm Fellowship, NSF NRT Award No. 2152260 and Ellis Fund administered by the University of California, Davis.

### v2

@@ 5 Methods and data properties
With sEMG articulations represented on the Riemannian manifold, we can apply methods described in Huang and Van Gool (2017) for SPD matrix learning to decode different articulations. We describe the architecture of the neural network in figure 14 and briefly recapitulate the core formulations in Huang and Van Gool (2017) here. The neural network architecture for learning discriminative SPD matrix representations is made of three layer types. First, a layer is defined by $\mathcal{E}_{k}=W_{k}^{T}\mathcal{E}_{k-1}W_{k}$, where an SPD edge matrix $\mathcal{E}_{k-1}$ of dimensions $c_{k-1}\times c_{k-1}$ is input to $k-$th layer. $W_{k}$ is of dimensions $c_{k-1}\times c_{k}$ giving rise to $\mathcal{E}_{k}$ of dimensions $c_{k}\times c_{k}$. $W_{k}$ is constrained to be a full-rank semi-orthogonal matrix on a compact Stiefel manifold such that $W_{k}^{T}W_{k}=I$. Second, a non-linear layer is defined by $\mathcal{E}_{k}$ = $U_{k-1}\max(\epsilon I,\Sigma_{k-1})U_{k-1}^{T}$. Third, a layer to map SPD matrices from the manifold space to its tangent space (so that the Euclidean operations can be applied) is defined by $\mathcal{E}_{k}=U_{k-1}\log(\Sigma_{k-1})U_{k-1}^{T}$. $U_{k-1}$, $\Sigma_{k-1}$ are obtained by eigendecomposition of matrix $\mathcal{E}_{k-1}$ and $\epsilon$ is a small constant $>0$.
@@ 6.1 Experiment 1
Word and phonetic articulations in table LABEL:tab:experimentPart1 are produced in 4 different sessions. Subjects articulated phonemes and words in an audible manner in the first 2 sessions and in a silent manner in the next 2 sessions. Within each session, subjects articulated a given word or phoneme 5 times. Therefore, we have 10 instances each of audible and silent articulations for all words and phonemes. We use first 3 instances from each session for training and last two instances for testing the models. Therefore, we have 6 instances of each word or phoneme for training and 4 instances for testing. Phonetic articulation data of subject 11 was corrupted during acquisition and is unavailable. For MDM and architecture in figure 14, edge matrices are created using the time-context window of 1.5s (the entire duration of articulation). For GRU architecture in figure 15, edge matrices are created with a sliding window of context size of 150ms and step size of 30ms. Results for the experiments in table LABEL:tab:experimentPart1 are given in tables 6, 4, and 5. The presented accuracy values are averaged over 10 random seeds (the model in figure 14 is trained for 1000 epochs and the model in figure 15 is trained for 150 epochs. Best accuracy on the test set in reported). Model in figure 14 contains about 5000 to 11000 parameters (depending on the size of the linear classification head) and the model in figure 15 has approximately 150000 parameters.
@@ Acknowledgments
This work was supported by awards to Lee M. Miller from: Accenture, through the Accenture Labs Digital Experiences group; CITRIS and the Banatao Institute at the University of California; the University of California Davis School of Medicine (Cultivating Team Science Award); the University of California Davis Academic Senate; a UC Davis Science Translation and Innovative Research (STAIR) Grant; and the Child Family Fund for the Center for Mind and Brain.
Harshavardhana T. Gowda is supported by Neuralstorm Fellowship, NSF NRT Award No. 2152260 and Ellis Fund administered by the University of California, Davis.
@@
Author contributions
  • Harshavardhana T. Gowda: Mathematical formulation, concepts development, data analysis, experiment design, data collection software design, data collection, manuscript preparation.
  • Zachary D. McNaughton: Data collection.
  • Lee M. Miller: Concepts development and
    manuscript preparation.